i am currently doing a section on this in my class called sloving linear equations by graphing and i dont understand how u do it can u help me?
Answer by Theo(13342)

you graph each equation separately.

the intersection of the graph of each equation is the solution that is common to all equations.

if in 2 dimensions, your equations would be something like:

5x + 6y = 3
3x - 2y = 7

you graph these equations by solving for y in each equation.

in the first equation of 5x + 6y = 3, you solve for y to get y = (-5/6)*x + (3/6)

in the second equation of 3x - 2y = 7, you solve for y to get y = (3/2)*x - (7/2)

now you graph both equation. the graph will look like this:



your answer will be the intersection of these lines.

that answer will be somewhere around x = 1.75 and y = -1.

to confirm, you need to solve the equations algebraically because graphing is difficult to pinpoint the solution exactly.

it allows you to get an idea where the actual solution will lie.

if you graph with the proper scale and you use graph paper, then the intersection will show up better than what I just showed you.

I'll solve these equations algebraically to show you where the actual solution lies.

the equations are:

5x + 6y = 3
3x - 2y = 7
multiply the second equation by 3 to get:

5x + 6y = 3
9x - 6y = 21

add the 2 equations together to get:

14x = 24

divide both sides of this equation by 14 to get x = 24/14 = 1.714285714

use that value of x to solve for y in the first equation to get y = -.928571429

the actual answer is:

x = 1.714285714 and y = -.928571429

as you can see from the graph, it would be impossible for you to spot that just by looking at the graph.

you can get the general vicinity, but you will not be able to get the exact answer.

solving by graphing is good to let you know if you even have a solution and where the solution will be near but it is not good to get the exact solution just by looking at the graph.